OpenStudy (anonymous):

What is the simplest form of the expression? Square root 2 minus 6 over 2 plus 6

OpenStudy (anonymous):

\[\frac{\sqrt2-6}{2+6}\]??

OpenStudy (anonymous):

my guess is it is actually \[\frac{\sqrt2-6}{\sqrt2+6}\]

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{6}\div \sqrt{2}+\sqrt{6}\]

OpenStudy (anonymous):

\[\frac{\sqrt2-\sqrt6}{\sqrt2+\sqrt6}\]

OpenStudy (anonymous):

I got \[8-4\sqrt{3}\div8\] is that correct

OpenStudy (anonymous):

multiply by the conjugate of the denominator start with \[\frac{\sqrt2-\sqrt6}{\sqrt2+\sqrt6}\times \frac{\sqrt2-\sqrt6}{\sqrt2-\sqrt6}\]

OpenStudy (anonymous):

seem unlikely the denominator will be \(2-6=-4\)

OpenStudy (anonymous):

would it not be 2+6 on never mind i see my error

OpenStudy (anonymous):

your numerator is right

OpenStudy (anonymous):

but you can cancel when you get the \(-4\) in the denominator

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[\frac{8-4\sqrt3}{-4}\] cancel the \(-4\)

OpenStudy (anonymous):

i.e. divide each term in the numerator by \(-4\)

OpenStudy (anonymous):

\[-2+\sqrt{3}\]

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

ok that makes since thank you

OpenStudy (anonymous):

yw